OpenStudy (anonymous):

A company with loud machinery needs to cut its sound intensity to 37% of its original level. By how many decibels would the loudness be reduced? Use the formula L = 10 Log I over I 0

OpenStudy (anonymous):

my guess is to set \[10\log(\frac{I}{I_0})=.37\]and solve for \(I\)

OpenStudy (anonymous):

i think it is best to ignore the \(I_0\) and solve via \[\log(I)=.037\] \[I=10^{.037}\]

OpenStudy (anonymous):

but i could be wrong

OpenStudy (anonymous):

@SolomonZelman sorry i was in jrotc and i had a teacher help me and and we came up with the equation L = 10 Log . 37 , close to yours. so now my question is do i add log to the equation when i plug it into my claculator or no ?

OpenStudy (anonymous):

* i was in jrotc . I had a teacher help me. sorry that was grammatically incorrect

OpenStudy (solomonzelman):

What is your formula that you are dealing with, can you type it in equations, so that I know what exactly you are talking about?